Consider a large one economy where:

Cd= 1+.9(Y-T)-150r

Id=80-200r

y=250 T=40 G=50 NFP=0

If the saving schedule for the rest of the world is Sd=15+516.6r and its investment schedule is Id=120-300r, then:

a) What is the equilibrium world interest rate?

b) What is the current account for the large open economy?

c) What is the financial account for the large open economy?
Answer #1

a)

At equilibrium,

Y = C+I+G

250 = 1+0.9(250-40)-150r + 80-200r + 50

250 = 320-350r

350r = 70

R* = 0.2 or 20%

b)

Current account = S-I

Current account = (15+516.6r) – (120-300r)

Current account = -105+216.6r

Current account = -105+216.6(0.2)

Current account = -$61.68

c)

Current account + Financial account = 0

-61.68 + financial account = 0

Financial account = $61.68